OpenStudy (anonymous):

A geophysicist has drilled a hole through the earth, from the north pole to the south pole. When she is at the south pole she jumps in to the hole and a some time later she comes up at the south pole again. (there is no air-resistence in the hole) a) How many seconds/minutes/ hours later does she come up again? b) How big is her maxvelocity? c) how big is her max acceleration? I know it will become like a harmonic oscillator but I dont know how to calculate it. thanks in advance.

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

within the earth the gravity distribution changes from the inverse square law to one were g is zero at the dead centre and increases linearly to g at the surface. so it would be just like being attached to a spring-like force with constant k where k = g/R where R is earth radius, with end of the relaxed spring being centred at earth's centre. if x is distance from dead centre you can begin to sub in directly to the standard shm equation x = A cost ( wt + delta) where these all have the usual meanings, delta being phase shift which you might wish to use here if x = R (ie A in the shm equation) at t = 0. in order for the person to re-appear at the S pole, they would have to go through 1 full cycle, is S to N to S pole. you could also do it by deriving the shm equations from force (but why bother), or by comparing energy levels again by analogy to the spring system, but i'd personally just have a crack at this the obvious way.

OpenStudy (irishboy123):

here is a really good exploration of the idea, with pictures and links that go beyond the simple idea, including a derivation of the gravity profile within the earth..... http://hyperphysics.phy-astr.gsu.edu/hbase/mechanics/earthole.html

OpenStudy (anonymous):

thanks @IrishBoy123

OpenStudy (anonymous):

In that link why is the period of oscillation equation = \[T=2\pi*\sqrt{R_e/g}\] how can m=R_earth and g=k? @IrishBoy123

OpenStudy (irishboy123):

ah, nice spot!! in simple horizontal (gravity independent) spring shm, we see m x'' + kx = 0. kx is measured in Newtons. this leads to w = root (k/m), with k = spring constant, m = mass here: m x'' + (mg/R) x = 0, where (mg/R) x is measured in Newtons in this case, the m's cancel. which makes complete sense from the physical reality. the mass of the person jumping in should be irrelevant. i think this was my error in my first post. so thank you. to summarise, the equivalent of k in this example, lets call it k' is k'= mg/R. thus, w = root ( k'/m) = root (g/R) i hope that helps...

OpenStudy (anonymous):

yeah thanks, but the maximum velocity? how do I derive that? and the max acceleration must be gravity at the earths surface? cause its just gets smaller the closer I get to earthcentre? @IrishBoy123

OpenStudy (irishboy123):

have you gone with x = R cost (wt - π/2) ?????? where: R = rad earth & w = root (g/R) if so x' = - wR sint(wt - π/2) x'' = - w^2 R cos(wt - π/2) cos and sin oscillate between 1 and -1 so: max/min vel = [+ or -] wR = root (g/R) * R = root (gR) maxmin accel = [+ or -] w^2 R = (g/R) * R = g

OpenStudy (anonymous):

No I dont think my math is that good to use derivate it like that. I did like this: at the earth surface \[F=mg\] and \[F_s=kR\] \[F_s=F >mg=kR>k=mg/R\] and a harmonic oscillator have the equations \[\omega=\sqrt{k/m}\] and \[T=2\pi/\omega\] So I first solved angular frequency: \[\omega=\sqrt{mg/R/m}>\omega=\sqrt{g/R}>\omega=0.00125\] then the frequancy in the Period equation. \[T=2\pi/\omega>2\pi/0.00125>5060 s \approx 84 minutes\] then for maximum speed, \[F=mg.F=mv^2/r>F=F> mg=mv^2/R_j>v=\sqrt{g*R_j}=7909,7m/s\] and maximum acceleration is 9.82m/s^2 Do you think i will get a passing grade for this assignment? @irishboy123

OpenStudy (anonymous):

@perl

OpenStudy (perl):

it can be shown that as you move inside the Earth, the Force acting on you will reduce linearly and is proportional to radius. I think you had that F = k*r

OpenStudy (irishboy123):

@pate16 i am happy to go through your approach. give me a bit so i can sit down and look. i still think that from a physics perspective you should stick to well established shm principles, and that from a cursory glance your maths looks well up to what is needed to apply the basic, ephemeral physics shm paradigm. i'll get back to you but perl might get you there in meantime. got to put my kids to bed.

OpenStudy (perl):

(Assume that R_e is the earth's radius) We know that for simple harmonic motion $$ \Large { T= 2\pi \sqrt{ \frac {m}{k}} \\ \text{But } ~F = k\cdot R_e ~~\to k = \frac {F}{R_e} \\ \therefore\\ T= 2\pi \sqrt{ \frac {m}{\frac {F}{R_e}}} T= 2\pi \sqrt{ \frac {m}{\frac {mg}{R_e}}} = 2\pi \sqrt{ \frac {R_e}{g}} } $$

OpenStudy (perl):

$$ \Large{ R_e = 6 378 100~ m\\ g = 9.81~ m/s^2\\ T = 2\pi \sqrt{ \frac{6 378 100}{9.81}} \approx5068.8856 ~sec \\ 5068.8856 ~sec = 84.5~ min }$$

OpenStudy (perl):

I don't quite follow your reasoning for finding maximum velocity

OpenStudy (anonymous):

No im not so sure of that either.. I saw it like a centripetal force but thats not quite true, is it?

OpenStudy (perl):

F= mv^2/r is used for uniform speed circular motion

OpenStudy (perl):

but the answer comes out approximately correct, so maybe you can use it

OpenStudy (anonymous):

really? but how would you do it?

OpenStudy (perl):

that approach would be used to find the speed of orbit for a satellite under the influence of gravity

OpenStudy (perl):

but in our problem we are doing linear motion, but I would have to ask @IrishBoy123 if this approach works. F = mv^2/r F =mg mg = mv^2 / r ---> v = sqrt( gr) maybe its a coincidence that it is the maximum velocity another way to approach this would be to make a velocity equation at time t

OpenStudy (perl):

I just checked this website, it is definitely correct. The maximum velocity is sqrt( R_e * g) http://www.quora.com/What-will-happen-when-a-ball-is-dropped-in-a-tunnel-that-passes-through-the-center-of-the-Earth

OpenStudy (anonymous):

oh nice haha, But could you show how you would do?

OpenStudy (perl):

here are some formulas http://www.schoolphysics.co.uk/age16-19/General/text/Formulae_list/index.html

OpenStudy (perl):

maximum velocity occurs when the derivative of velocity is zero , so when acceleration is equal to zero

OpenStudy (perl):

we can use the equations for simple harmonic motion $$ \Large{ \\ displacement = r \sin ( \omega t ) \\ velocity = \omega ~r \cos(\omega t) \\ acceleration = -\omega^2r\sin(\omega t) }$$

OpenStudy (perl):

Let me give you an intuitive explanation, there are more general approaches that deal with it more generally.

OpenStudy (perl):